a disc of 8cm radius rotate horizontally with angular velocity 120 revolution per second. Suddenly a small particle drop onto the edge of disc. The angular velocity of the disc reduce to 85 revolution per second. Calculate the mass of small particles if moment inertia of the disc is 6x10^-4 kgm²


1
Expert's answer
2021-10-02T14:55:43-0400

"I_1\omega_1=(I_1+mr^2)\cdot\omega _2\to m=\frac{I_1\omega_1-I_1\omega_2}{r^2\omega_2}="


"=\frac{6\cdot 10^{-4}\cdot 754-6\cdot 10^{-4}\cdot 534}{0.08^2\cdot 534}=0.0386\ (kg)=38.6\ (g)" . Answer